OpenStudy (ilovelife):

What is the solution set of |x – 6| = 8 ? So, do i do x + 6 = 8 and then x - 6 = 8, and that's how i find my answer ?

jimthompson5910 (jim_thompson5910):

first one should be x-6 = 8 and the second part should be x-6 = -8

jimthompson5910 (jim_thompson5910):

solve each to find your two answers

OpenStudy (accessdenied):

The absolute value means that, regardless of whether 'x-6' is positive or negative, it always \(becomes\) positive. |x - 6| = 8 <--- Regardless of whether 'x-6' is +8 or -8, the absolute value of it is always +8. So, it breaks into: x - 6 = 8 or x - 6 = -8.